# Float Self-Tagging

### v1

@@ 3.3. Memory Allocation
Memory usage is measured by counting the number of heap-allocated bytes across the lifetime of each benchmark. The original and self-tagging Bigloo versions are then compared to assess the efficiency of self-tagging at unboxing floats and its impact on benchmarks that make limited use of floats. Results are shown in Figure 5.
Figure 5. Heap allocations comparison between the original Bigloo version and the self-tagging version. The y-axis shows heap allocations relative to the original version. 1.0 indicates no change, while 0.0 indicates that all allocations were removed.
On all benchmarks extensively using floats, memory allocation decreases. On mbrot, nucleic, simplex and sum1, allocations decreases by about 25%, which is consistent with the proportion of self-tagged floats ($1/4$). The fft benchmark has only 17% fewer allocations due to predominantly using the float 0.0, which the original Bigloo version preallocates. For fibfp, pnpoly, ray and sumfp there is very little allocation of floats when using self-tagging because those programs mostly use “simple” floats that are integer valued floats $\leq 2^{51}$ scaled by a small power of two, such as 0.75, 2.25, or 42.0.
On all benchmarks that use few floats, memory usage increases by less than 1%. This tenous increase stems from the use of generic pointers for types that were assigned specific tags in the original version.
@@ 3.4. Execution Time
Figure 6 shows the execution time comparison between both implementations. Execution time refers to elapsed real time measured with perf-stat. Each benchmark is parameterized for a run to last for at least 10 seconds, and is repeated 5 times. All timing results are the average execution time of these runs.
Figure 6. Execution time comparison between the original Bigloo version and the self-tagging version. The y-axis shows execution time relative to the original version. 1.0 indicates no change. Lower means faster execution than the original version.
@@
This unpredictability is explainable, and fixable. Figure 7 shows that, despite reducing float allocations in mbrot and nucleic, self-tagging also causes about 6 times more missed branch predictions. This stems from the fact that tags occupy the low bits of the mantissa. In that position, the smallest variations cause a float’s representation to switch between tagged value and tagged pointer in a hard to predict way. This is beyond the capabilites of the branch predictor and execution suffers costly misprediction penalties (Bryant and O’Hallaron, 2016).
Figure 7. Missed branch predictions comparison between the original Bigloo version and the self-tagging version. The log-scaled y-axis shows relative number of missed branch predictions compared the original version. 1.0 indicates no change, lower means fewer missed predictions and higher means more missed predictions.
@@ 4. Improved Predictability and Coverage
This section addresses the issue uncovered by the previous experiment and presents an improved, performant implementation of float self-tagging.
The previous experiment shows that while self-tagging decreases the number of heap-allocated floats it does not consistently improve performance due to failed branch predictions. To fix this issue, the tag system can be modified to bias it in favor of common floats by superimposing the 3 tag bits with the most significant bits of the exponent, as shown in Figure 8. Hence, all floats whose exponent falls into a given range, which is determined by the tag, are unboxed. Since tags must be placed on low bits, floats are encoded to tagged values by a 60-bit rotation to the right (or 4-bit to the left). Decoding is achieved with the inverse rotation.
@@
Given a tag, the floating-point intervals it covers can be computed as follows. Assuming a 0 sign bit, the lowest and highest bounds are reached when all eight exponent low bits and all mantissa bits are set to 0 and 1 respectively. There are two exceptions when the tag is either 000 or 111. The former represents subnormal numbers when all exponent bits are set to 0, in which case the lowest bound is 0.0. In the case of the tag 111, when all exponent bits are set to 1, the floating-point representation is that of Infinity or NaN. Since the sign bit is not part of the tag, each tag also captures a corresponding negative interval.
Exponent Value
Tag range range fft fibfp mbrot nucleic pnpoly ray simplex sumfp sum1
0.0 94% 13% 0% 0% 16% 13% 11% 0% 0%
000 0..255 $2^{-1074}$..$2^{-767}$
001 256..511 $2^{-767}$..$2^{-511}$
010 512..767 $2^{-511}$..$2^{-255}$
011 768..1023 $2^{-255}$..2222 6% 33% 96% 70% 60% 18% 62% 0% 0%
100 1024..1279 $2$..$2^{257}$ 0% 54% 4% 30% 24% 69% 27% 100% 100%
101 1280..1535 $2^{257}$..$2^{513}$
110 1536..1791 $2^{513}$..$2^{769}$
$2^{769}$..$2^{1024}$
111 1792..2047 Infinity/NaN - - - - - - - - -
(a) The positive value ranges captured by each tag and the proportion of all floats generated by the R7RS float-intensive benchmarks. The lowest and highest exponents are exponents before applying the bias ($-1023$ for 64-bits floats). Values in the interval formed by the lower bound (included) and upper bound (excluded) can be unboxed by self-tagging for the given tag. Intervals of interest are highlighted. Empty entries indicate no float in this range was generated, whereas 0% means that very few floats in this range were generated (less than 0.5%).
(b) Logarithmic scale number line of float values, excluding NaN values and infinities. Intervals of interest are highlighted with their corresponding color from Figure 9(a).
Figure 9. Floating point interval captured by each tag.
Figure 9(a) shows the positive interval captured by each tag. Three tags stand out as capturing worthwhile intervals. The tags 011 and 100 capture the range of floats from $2^{-255}$ to $2^{257}$, or about $1.7\times 10^{-77}$ to $2.3\times 10^{77}$. The tag 000 captures the values $\pm$0.0 and, incidentally, the floats between $-2^{-767}$ and $2^{-767}$. Figure 9(b) shows the intervals captured by these tags on a number line.
@@
Contrarily to the implementation of self-tagging in Section 3, these tags capture intervals of contiguous values. This contiguity within intervals prevents small variations from causing switches from a tagged value to a tagged pointer representation.
In fact, although the tagged floats cover only $3/8$ of all floats, it is reasonable to expect that many programs will use floats in these intervals with very high probability, which benefits branch prediction. The right part of Figure 9(a) confirms this expectation. It presents the distribution of floats that are generated in each float-intensive R7RS benchmark. Values are aggregated by tag, with the exception of $\pm$0.0 and Infinity/NaN, which are counted separately. The tags 000, 011 and 100 cover all floats in these benchmarks. Moreover, among the values captured by 000, only $\pm$0.0 are used.
This permits dropping the tag used for the tagged pointer representation of the remaining floats due to their expected rarity. They are instead represented with generic pointers with a header type. Whether expending the tag 000 to capture 0.0 and -0.0 is a desirable trade-off is discussed in the next section.
4.1. A Second, Improved Experiment
The experimental setup from Section 3 is replicated to test float self-tagging using high bits of the exponent. Two variants are compared to the original version of Bigloo that implements standard tagged objects.
  • self-tagging (3 tags) uses the tags 000, 011 and 100, thus unboxing floats in the intervals from Figure 9(b). Remaining floats are represented by generic pointers with a header type.
  • self-tagging (2 tags) tests the impact of expending the tag 000 for self-tagging. It only uses the tags 011 and 100 for float self-tagging and reclaims the tag 001 to represent all remaining floats as tagged pointers. The floats $\pm$0.0 are preallocated.
Figure 10 shows the memory allocations of the improved variants. Both variants allocate fewer floats on all float-intensive benchmarks. In fact, both versions have been verified to allocate no float at all (see Figure 9(a)) with the only exceptions being the preallocated $\pm$0.0 by the second version. The remaining heap usage is for non-float objects. Similarly to the previous experiment (Figure 5), memory usage of other benchmarks remains mostly unaffected.
Figure 10. Heap allocations comparison between both versions with self-tagging on exponent bits. The 3 tags version uses tags 000, 011 and 100. The 2 tags version uses only 011 and 100, but preallocates 0.0 and -0.0. The y-axis shows heap allocations relative to the original Bigloo version, which uses tagged pointers for all floats. 1.0 indicates no change, while 0.0 indicates that all allocations were removed.
Figure 11 shows faster execution times for both variants of self-tagging. All float-intensive benchmarks execute faster than the original Bigloo implementation, with the exception of fft on the version that does not use the tag 000. In particular, the mbrot and nucleic benchmarks are both faster, confirming that self-tagging on exponent bits eliminates the branch misprediction penalty incurred by self-tagging on mantissa low bits (see Section 3). Despite small individual variations, other benchmarks remain generally unaffected.
The version using three tags for self-tagging, including the tag 000, is 2.3$\times$ faster on average than the original Bigloo version. This is significantly higher than the 1.7$\times$ acceleration without using the tag 000. In particular, the benchmark fft, which predominantly uses the values $\pm$0.0, is slower in the second version despite zeros being preallocated. Hence, it is better performance-wise to include the tag 000.
Figure 11. Execution time comparison between both versions with self-tagging on exponent bits. The 3 tags version uses tags 000, 011 and 100. The 2 tags version uses only 011 and 100, but preallocates 0.0 and -0.0. The y-axis shows execution time relative to the original version. 1.0 indicates no change, lower means faster execution than the original Bigloo version, which uses tagged pointers.
4.2. Application to JavaScript
This section experiments with the application of float self-tagging in JavaScript, which is a prime candidate for self-tagging because it has no integer type. Its specification defines numbers as double-precision floats (International, 2015). To isolate the impact of the representation of float, three versions of the Hopc JavaScript compiler are tested, where each version differs only by its representation of floats, which is either tagged pointers, NaN-tagging or self-tagging.
Hopc is an ahead-of-time JavaScript compiler that uses Scheme as its compilation backend and relies on the underlying Scheme compiler for optimizations (Serrano, 2021). This permits reusing Bigloo’s runtime to test float self-tagging in the context of JavaScript. Bigloo can also be configured to use NaN-tagging. This configuration is scarcely used for Scheme because of its negative impact on other types, but is applicable to compilation of JavaScript.
Figure 12 compares Bigloo’s NaN-tagging performance to that of float self-tagging with three tags from Section 4.1. On float-intensive benchmarks, self-tagging generally fares slightly better than NaN-tagging. Moreover, on non-float benchmarks the float-bias of NaN-tagging makes it slower than the original implementation on average, while the self-tagging version remains generally unaffected.
Figure 12. Execution time comparison between Bigloo with self-tagging on exponent bits and Bigloo with NaN-tagging. The y-axis shows execution time relative to the original Bigloo version, which uses tagged pointers. 1.0 indicates no change, lower means faster execution than the original version.
Figure 13 presents the results of porting Bigloo’s NaN-tagging and self-tagging implementations to JavaScript. Performance are measured using the Hop benchmark suite, which gathers benchmarks from the popular JetStream, Octane and Sunspider JavaScript benchmark suites (Serrano, 2024). Once again, it is useful to categorize benchmarks between float-intensive benchmarks (almabench, n-body, navier-stokes, and raytrace), and benchmarks that use few or no floats.
The Hopc NaN-tagging and self-tagging implementations have comparable performance on float-intensive benchmarks. Execution time with NaN-tagging is faster on almabench and n-body, but slower on raytrace, and equivalent on navier-stokes. However, on other benchmarks the self-tagging implementation generally fares better due to its lower overhead when handling non-float objects.
Figure 13. Execution time comparison between Hopc with self-tagging on exponent bits and Hopc with NaN-tagging. The y-axis shows execution time relative to the original Hopc version, which uses tagged pointers. 1.0 indicates no change, lower means faster execution than the original version.
@@ 5. Implementation
This section discusses the implementation of the float-related operations of the dynamic language. The type f64 will refer to IEEE754 64-bit doubles and the type i64 will refer to 64-bit integers. Moreover, the type object will refer to the values of the language (the sum type of floats, integers, booleans, and other objects), which is an i64. To better understand the low-level implications of the implementation, assembly code required on x86-64 is shown to illustrate the details of the implementation on that architecture.
5.1. Tag Testing
Testing the low 3-bit tag of an i64 value is a basic need for float-related operations and dynamic type checks. Checking that the 3 low bits are equal to a specific tag can be achieved by masking those bits followed by a comparison. Assuming the i64 value is in the x86 64-bit register rax, then the following sequence will branch to label matching_tag when the 3 low bits are equal to tag:
and al,#3
cmp al,#tag
jz matching_tag
Note that by using al only the lower 8 bits of rax are accessed, which leads to smaller constants and instruction encodings. Unfortunately, this sequence modifies rax so if the value is needed after the test, as often will be the case, then an additional instruction and register are required to keep a copy. In the special case of tag=0, the two first instructions can be replaced with “test al,#7”, which also eliminates the need for a copy because it does not modify rax.
Because a tagged float will have one of a set of tags, it is useful to check this more efficiently than a sequence of comparisons. This can be done with the bit-test instruction “bt n,i” that reads the bit at index i of n and puts it in the carry flag. The bt instruction comes in 16, 32, and 64 bit variants, but not 8 bit. As an example, testing that the least 3 low bits of rax are either 000 (0), 011 (3), or 100 (4), can be done with:
mov bx,#0x1919 # set of tags

### v2

@@ 3.3. Memory Allocation
(a) Relative memory allocation
(b) Relative execution time
(c) Relative missed branch prediction
Figure 5. Heap allocations, relative execution time and missed branch predictions comparison between the original Bigloo version and the self-tagging version. The y-axis shows each measure relative to the original version. 1.0 indicates no change. For all metrics, lower is better.
Memory usage is measured by counting the number of heap-allocated bytes across the lifetime of each benchmark. The original and self-tagging Bigloo versions are then compared to assess the efficiency of self-tagging at unboxing floats and its impact on benchmarks that make limited use of floats. Results are shown in Figure 5(a).
On all benchmarks extensively using floats, memory allocation decreases. On mbrot, nucleic, simplex and sum1, allocations decrease by about 25%, which is consistent with the proportion of self-tagged floats ($1/4$). The fft benchmark has only 17% fewer allocations due to predominantly using the float 0.0, which the original Bigloo version preallocates. For fibfp, pnpoly, ray and sumfp there is very little allocation of floats when using self-tagging because those programs mostly use “simple” floats that are integer valued floats $\leq 2^{51}$ scaled by a small power of two, such as 0.75, 2.25, or 42.0.
On all benchmarks that use few floats, memory usage increases by less than 1%. This tiny increase stems from the use of generic pointers for types that were assigned specific tags in the original version.
@@ 3.4. Execution Time
Figure 5(b) shows the execution time comparison between both implementations. Execution time refers to elapsed real time measured with perf-stat. Each benchmark is parameterized for a run to last for at least 10 seconds, and is repeated 5 times. All timing results are the average execution time of these runs.
@@
This unpredictability is explainable, and fixable. Figure 5(c) shows that, despite reducing float allocations in mbrot and nucleic, self-tagging also causes about 4 times more missed branch predictions. This stems from the fact that tags occupy the low bits of the mantissa. In that position, the smallest variations cause a float’s representation to switch between tagged value and tagged pointer in a hard to predict way. This is beyond the capabilites of the branch predictor and execution suffers costly misprediction penalties (Bryant and O’Hallaron, 2016).
5 most 5 most
signif. signif.
expo. expo. Value
bits bits + 1 range fft fibfp mbrot nucleic pnpoly ray simplex sumfp sum1
0.0 94% 13% 0% 0% 16% 13% 11% 0% 0%
00000 00001 5e-324 .. 2.1e-289 - - - - - - - - -
00001 00010 2.1e-289 .. 3.8e-270 - - - - - - - - -
00010 00011 3.8e-270 .. 7e-251 - - - - - - - - -
00011 00100 7e-251 .. 1.3e-231 - - - - - - - - -
00100 00101 1.3e-231 .. 2.4e-212 - - - - - - - - -
00101 00110 2.4e-212 .. 4.4e-193 - - - - - - - - -
00110 00111 4.4e-193 .. 8.1e-174 - - - - - - - - -
00111 01000 8.1e-174 .. 1.5e-154 - - - - - - - - -
01000 01001 1.5e-154 .. 2.8e-135 - - - - - - - - -
01001 01010 2.8e-135 .. 5.1e-116 - - - - - - - - -
01010 01011 5.1e-116 .. 9.4e-97 - - - - - - - - -
01011 01100 9.4e-97 .. 1.7e-77 - - - - - - - - -
01100 01101 1.7e-77 .. 3.2e-58 - - - - - - - - -
01101 01110 3.2e-58 .. 5.9e-39 - - - - - - - - -
01110 01111 5.9e-39 .. 1.1e-19 - - - - - - - - -
01111 10000 1.1e-19 .. 2 6% 33% 96% 70% 60% 18% 62% 0% 0%
10000 10001 2 .. 3.7e19 0% 54% 4% 30% 24% 69% 27% 100% 100%
10001 10010 3.7e19 .. 6.8e38 - - - - - - - - -
10010 10011 6.8e38 .. 1.3e58 - - - - - - - - -
10011 10100 1.3e58 .. 2.3e77 - - - - - - - - -
10100 10101 2.3e77 .. 4.3e96 - - - - - - - - -
10101 10110 4.3e96 .. 7.9e115 - - - - - - - - -
10110 10111 7.9e115 .. 1.5e135 - - - - - - - - -
10111 11000 1.5e135 .. 2.7e154 - - - - - - - - -
11000 11001 2.7e154 .. 4.9e173 - - - - - - - - -
11001 11010 4.9e173 .. 9.1e192 - - - - - - - - -
11010 11011 9.1e192 .. 1.7e212 - - - - - - - - -
11011 11100 1.7e212 .. 3.1e231 - - - - - - - - -
11100 11101 3.1e231 .. 5.7e250 - - - - - - - - -
11101 11110 5.7e250 .. 1.1e270 - - - - - - - - -
11110 11111 1.1e270 .. 1.9e289 - - - - - - - - -
1.9e289 .. 1.8e308 - - - - - - - - -
11111 00000 Infinity/NaN - - - - - - - - -
Figure 6. The positive value ranges captured by each combination of the 5 most significant exponent bits and the proportion of all floats computed by the float-intensive programs in the R7RS benchmark suite. Intervals of interest are highlighted with a background color. Empty entries indicate no float in this range was computed, whereas 0% means that very few floats in this range were generated (less than 0.5%).
@@ 4. Improved Predictability and Coverage
This section addresses the issue uncovered by the previous experiment and presents an improved, performant variant of float self-tagging.
The previous experiment shows that while self-tagging decreases the number of heap-allocated floats it does not consistently improve performance due to failed branch predictions. To address this issue, self-tagging is used for the most common floats encountered in practice to maximize the likelihood that a float will be self-tagged, thus improving branch prediction and decreasing heap-allocated floats.
But which floats are most common in programs? To answer this question the float-intensive programs of the R7RS benchmark suite are instrumented to create a profile of all the floats computed by those programs. The results are shown in Figure 6. The table contains 32 rows, one for each combination of the 5 most significant bits of the float’s 11-bit exponent field, which is indicated in the first column. For now ignore the second column. Each of the right columns indicates what proportion of computed floats have a specific combination of most significant exponent bits. The computed floats are clearly clustered around 2, specifically in the range $1.1\times 10^{-19}$. $3.7\times 10^{19}$, and the value 0.0 also occurs frequently for some programs. It is clear that other scientific programs may exhibit profiles that use other ranges of floats. In particular, Infinity and NaN may show up more often in some types of computations. It is apparent that many programs will operate with floats that are predominantly those around 2 on a logarithmic scale, and also 0.0, Infinity, and NaN.
4.1. Self-Tagging with 3 Tags and 4 Tags
@@
The float profile of Figure 6 has the nice property that all floats computed have the 3 most significant bits of the exponent field that are either 000 (green rows), 011 (blue rows), or 100 (orange rows). Moreover, among the values captured by 000, only $\pm$0.0 are used.
These 3 exponent bits can be shifted using a bit rotation to align them with the lower 3 tag bits as shown in Figure 7. Thus a float is encoded to a tagged value by a 4-bit rotation to the left (or 60-bit to the right). A tagged value is decoded to a float with the inverse rotation.
@@
Contrarily to the implementation of self-tagging in Section 3, these tags capture intervals of contiguous values. This contiguity within intervals prevents small variations from causing switches from a self-tagged value to a heap-allocated float.
Test programs did not compute the float values $\pm$±Infinity and NaN, but those values would be assigned the tag 111 (gray rows). Assigning this fourth tag to self-tagged floats might be useful to avoid heap-allocated floats in programs frequently computing these values.
Given the expected rarity of the remaining floats there isn’t much of an advantage to represent them using tagged pointers. Instead a representation using a generic pointer with a header type is a better tradeoff as it allows using the tag that would have been used for some other, more frequently used type.
4.2. Self-Tagging with 2 Tags
As is shown in the float profile of Figure 6 the tag 000 is only used to capture the values $\pm$0.0. It is possible to avoid using the tag 000 for self-tagging by preallocating the values $\pm$0.0 using the generic pointer representation. When a float must be encoded as an object it is compared to zero, in which case one of the preallocated zeroes is returned, thus avoiding a new allocation. The same can be done for the values $\pm$±Infinity and NaN, which are expected to be the most frequent values tagged with 111. If this is done then only 2 tags are needed to cover the most frequent cases: 011 and 100.
Unfortunately, this approach may raise the number of failed branch predictions due to the test for the special values $\pm$0.0, $\pm$±Infinity, and NaN. Consequently it may be a more performant tradeoff to avoid testing for some of the special values which are instead handled with a heap-allocation. The best choice will depend on the context, for example the float profile of Figure 6 would suggest not testing for $\pm$±Infinity and NaN.
4.3. Self-Tagging with 1 Tag
Self-tagging can also be achieved with a single tag by a simple transformation of the most significant exponent bits. The 5 most significant bits of the exponent are added to 1 and the middle 3 bits are used as the tag (using a rotation 5 places to the left). This corresponds to the second column of Figure 6. Note that the tag 000 (bold pink rows) covers all the useful ranges, including $\pm$±Infinity and NaN that are not covered by the 2-tag and 3-tag variants.
It is noteworthy that there is nothing special about the tag 000. If for some external reason some other tag is more convenient it can easily be assigned to 1-tag self-tagged floats by adding $1+2\times tag$ rather than 1 to the 5 most significant exponent bits.
Putting all of this together, the self-tagged encoding of a float whose IEEE754 64-bit representation is the integer $n$ can be computed as $(n\oplus((1+2\times tag)<<58))<<_{rot}5$, where $\oplus$ is addition modulo $2^{64}$ and $<<_{rot}$ is the left rotation operator. The operations are reversed to perform the decoding of a self-tagged float to its IEEE754 64-bit representation.
This variant of self-tagging has the advantages of covering all the useful ranges of float values and using a single tag, which leaves more tags available for representing other objects. The only downside when compared to the 4-tag variant is that it covers proportionately fewer ranges of floats, so there is a risk of more heap-allocated floats.
4.4. Self-Tagging with a Different Exponent Bias
A small change in the exponent bias of the 1-tag variant can be used to get another self-tagging variant. If $2\times tag_{1}$ is added to the highest 5 bits and the middle 3 bits are examined, then when they are equal to $tag_{1}$ or $tag_{2}=tag_{1}-1$ they are self-tagged floats. This effectively doubles coverage by including both the light and bold pink rows of the second column of Figure 6. In other words the following (absolute value) ranges will be self-tagged: $0..3.8\times 10^{-270}\cup 5.9\times 10^{-39}..6.8\times 10^{38}\cup 1.1\times 10^{270}..Infinity/NaN$. To put this into perspective, this is a superset of the IEEE754 32-bit floats, so any floating point value that can be represented as a IEEE754 32-bit float can be self-tagged with this variant which uses 2 of the 8 3-bit tags. This is very likely a better use of 2 tags as the 2-tag variant of Section 4.2 does not cover the ranges containing 0, Infinity, and NaN which need special treatment that slows down execution.
When compared to the 1-tag variant, there will be no difference in heap allocations for programs with similar float profiles as Figure 6. However, as will be explained in the next section, it may yield more compact and faster code, in essence because the bias can be added after the rotation, and the need for a bias disappears when $tag_{1}$=000.
@@ 5. Implementation
This section discusses the implementation of the float-related operations of the dynamic language. The types f64 and f32 will refer to IEEE754 64-bit and 32-bit doubles and the types i64 and i32 will refer to 64-bit and 32-bit integers. Moreover, the type object will refer to the values of the language (the sum type of floats, integers, booleans, and other objects). To better understand the low-level implications of the implementation and give a sense of the execution time performance, assembly code for the Intel x86 architecture is shown to illustrate the details of the implementation on that popular architecture. Although this section focusses on a 64-bit implementation where the type object is an i64 and floats are f64, it is fairly straightforward to implement the operations similarly on a 32-bit implementation where the type object is an i32 and floats are f32.
Testing the low tag bits of an object value is a basic need for float-related operations to convert f64 to and from object and for dynamic type checks. In particular it is possible to determine if a f64 can be self-tagged by going through the encoding process and checking that the resulting tag is one of the tags (or the tag) appropriate for the chosen self-tagging variant.
5.1. Single Tag Testing
Checking that the N low bits are equal to a specific tag can be achieved by masking those bits followed by a comparison. Assuming the object value is in the x86 64-bit register rax, then the following sequence will branch to label tag_matches when the 3 low bits are equal to tag [^1] Intel assembler syntax is used throughout.:
and al, 7
cmp al, $tag$
jz tag_matches
Note that by using al only the lower 8 bits of rax are accessed, leading to smaller constants and instruction encodings. However, this sequence modifies rax so if the value is needed after the tag test, as often will be the case, then an additional instruction and register are required to keep a copy. In the special case of tag=0, the two first instructions can be replaced with “test al, 7”, which is a bitwise-and that tests if all 3 lowest bits of rax are zero without modifying rax. In the general case when $\neq$0 the test instruction can be used by combining it with a lea (Load Effective Address) instruction that copies to a temporary register the source register plus a constant that cancels tag (modulo 8):
lea ebx, [eax+(8-$tag$)]
test bl, 7
jz tag_matches
5.2. Multiple Tag Testing
When it is necessary to check if an object’s tag is one of a set of tags it is often possible to do this more efficiently than a sequence of single tag checks. This is particularly the case when there is leeway in the assignments of tags as is often the case.
[…]
lea ebx, [eax+(8-3)] # 011 -> 000 and 100 -> 001
test bl, 6
jz tag_matches_011_or_100
This technique can be extended when testing the 4 tags 000, 011, 100, and 111 of the 4-tag self-tagging approach:
lea ebx, [eax+(8-7)] # 000 -> 001, 011 -> 100, 100 -> 101, and 111 -> 000
test bl, 2
jz tag_matches_000_or_011_or_100_or_111
[…]
mov bx, 0x1919 # set of tags